OpenStudy (anonymous):

Simplify the expression sqrt(8x^3)+x sqrt (2)+4 sqrt(3x)

OpenStudy (anonymous):

Remember sqrt(a*b)=sqrt(a)*sqrt(b)

OpenStudy (mertsj):

\[\sqrt{8x^3}+x \sqrt{2}+4\sqrt{3x}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

@Mertsj its an x on the outside of the 8x^3 as well

OpenStudy (anonymous):

@Xavier that applies to this as well

OpenStudy (mertsj):

\[x \sqrt{8x^3}+x \sqrt{2}+4\sqrt{3x}=x \sqrt{4(2)x^2x}+x \sqrt{2}+4\sqrt{3x}\]

OpenStudy (mertsj):

\[x(2)x \sqrt{2x}+x \sqrt{2}+4\sqrt{3x}=2x^2\sqrt{2x}+x \sqrt{2}+4\sqrt{3x}\]

OpenStudy (anonymous):

o i see what you did ok